A1.30kg nano -satelite (a cube Salt )with a side length of 100.0mmis sitting on a workshop bench on the moon.

The acceleration due to gravity is 1.6 ms

calculate the pressure ?


1
Expert's answer
2022-06-30T09:54:11-0400

"p=\frac FS=\frac {mg}{a^2}=208~Pa."